Albert's father is 24 years old than Albert is. In 8 years, his father will be twice as old as Albert. How old is Albert now? How old is his father?

Found 2 solutions by macston, addingup:
Answer by macston(5194)

.
A=Albert's age; F=Father's age=A+24
.
F+8=2(A+8)
(A+24)+8=2A+16
A+32=2A+16
16=A
ANSWER 1: Albert is 16 years old.
.
F=A+24=16+24=40
ANSWER2: Father is 40 years old.
.
CHECK:
F+8=2(A+8)
40+8=2(16+8)
48=2(24)
48=48
.

Answer by addingup(3677)

f= A+24
A+8= (f+8)/2 Substitute for f:
A+8= ((A+24)+8)/2
2A+16= A+24+8
2A+16= A+32
A= 16
Albert is 16 and his father:
16+24= 40 In 8 years his father will be twice as old as Albert:
16+8= 24 and 40+8= 48, and of course 24*2=48. We have the correct answer.